Answers should be written in sentence form. Explanations must be detailed and well-thought out. Include examples wherever possible.
3. Explain why you might need to multiply each equation by a different number when solving a system of equations by ELIMINATION.
Please and thank you very much.

Answer by stanbon(75887)

. Explain why you might need to multiply each equation by a different number when solving a system of equations by ELIMINATION
--------------
Elimination requires the coefficients of a variable in two equations
are either the same or the negative on one another. If one is not
the multiple of the other, both must be multiplied by different
numbers so the products are equal or one is the negative of the other.
---------------
Example:
2x + 3y = 7
5x - 2y = 5
-----------------
You may chose to eliminate the x variable or the y variable.
To eliminate the x variable, Multiply the 1st equation by 5
and the 2nd equation by 2 to get:
-----
10x + 15y = 35
10x - 4y = 10
------------
Subtracting will eliminate the x terms and leave you with an
equation in y that can be solved.
---------------
Cheers,
Stan H.